OpenStudy (anonymous):

Help me out with this step by step please.... Write the equation of the line that is perpendicular to the line y = 3/2x + 4 and passes through the point (−6, 3).

OpenStudy (anonymous):

answers are: y = 3/2x - 1 y = 3/2x + 3 y = - 2/3x -1 y = -2/3x + 3

OpenStudy (anonymous):

please help

OpenStudy (bibby):

The slope of a perpendicular line has a negative reciprocal of the slope of the original line. The slope here is 3/2, right? What would the perpendicular lines slope be

OpenStudy (anonymous):

yes

OpenStudy (bibby):

Spoiler: that wasn't a yes/no question. The answer is a number

OpenStudy (anonymous):

What do you mean the answer is a number

OpenStudy (anonymous):

oh I get it

OpenStudy (anonymous):

It would be -2/3 ?

OpenStudy (bibby):

yep :D Now we can scratch off A and B

OpenStudy (anonymous):

yes we can

OpenStudy (anonymous):

now what?

OpenStudy (bibby):

Now we test C and D to see if they pass through the given point. Do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (bibby):

Alright. At this point we have our answers narrowed down to y = - 2/3x -1 y = -2/3x + 3 and the point (−6, 3). These are equations so we can either plug in 3 for the y value or -6 for the x value. Either way works. If you plug in -6 for x you should get a 3 from the equation Vice versa if you use 3 for y

OpenStudy (shamil98):

You don't have to plug in you have -2/3 = m and (-6,3) --> x,y plug them in and solve for b.. 3 = -2(-6)/3 + b

OpenStudy (shamil98):

plug them into slope-intercept form i should say y = mx +b

OpenStudy (bibby):

wat they already are in slope intercept form. my way should work too anyways 3 = -2/3x - 1 4 = -2/3x -6 = x

OpenStudy (shamil98):

your way you have to test each option , my way you can find it out without the answer choices lol. less tedious

OpenStudy (anonymous):

your both right just one way is easier then the other

OpenStudy (bibby):

I love you shamil

OpenStudy (shamil98):

m = -2/3 (x,y) --> (-6,3) y = mx + b 3 = -2(-6)/3 + b 3 = 4 +b -1 = b ily to man

OpenStudy (shamil98):

y = -2/3x - 1 and there u go.

OpenStudy (bibby):

Please give my associate a medal izzy

OpenStudy (anonymous):

how do i give medals?

OpenStudy (shamil98):

no give bib the medal i just interrupted

OpenStudy (shamil98):

click on best response next to bibby's name

OpenStudy (bibby):

bibby is a funny way of spelling shamil :*

OpenStudy (anonymous):

opps sorry clicked the wrong one

OpenStudy (bibby):

Glad we could help